OpenStudy (anonymous):

{(2/3)+3}/{3(2/3)-2} = 36666666666666666666666666666666666666666666666666666666666666667 However, whenever I input this in the website says it's wrong. Where did I make the mistake?

OpenStudy (zzr0ck3r):

\[\frac{\frac{2}{3}+3}{3\frac{2}{3}-2}=\frac{\frac{11}{3}}{0}\text{ is not a number}\]

OpenStudy (anonymous):

Good god, thank you!

OpenStudy (zzr0ck3r):

np, calculators do not like infinity:)